Solve the equation. Identify each as a conditional equation, an inconsistent equation, or an identity.
(4)x/4=x
2(0.5x+1.5)-3.5=3(0.5X+0.5)

Answer by user_dude2008(1861)

4x/4=x
x=x .... identity


2(0.5x+1.5)-3.5=3(0.5x+0.5)
x+3-3.5=1.5x+1.5
x-0.5=1.5x+1.5
-0.5x=2
x=-4 ... conditional